OpenStudy (anonymous):

How do you simplify -3 (rad24) +3 (rad6)-3 (rad18)?

OpenStudy (anonymous):

so it is -3 * 24 radians + 3 * 6 radians - 3 * 18 radians?

OpenStudy (campbell_st):

well look at it this was \[\sqrt{24} = \sqrt{4 \times 6} = \sqrt{4} \times \sqrt{6} = 2\sqrt{6}\] try to re-write each radical in terms of factors.... and 1 factor needs to me a square number... so using the example above \[-3\sqrt{24} = -3 \times 2 \sqrt{6} = -6\sqrt{6}\] leave the 2nd term alone and the last radical 18 = 9 x 2

OpenStudy (anonymous):